How to calculate iodine number given 0.11g sample fat? Solution contains 5 mL cyclohexane, 7 mL Wijs reagent, 7 mL 10% KI. Titrated with 0.0500M sodium thiosulfate, with 0.0248 L required to reach end point.
1
Expert's answer
2013-12-05T10:04:58-0500
Principle of Wijs method:
The vegetable oil sample is reacted with an excess of iodine monochloride
solution (Wijs reagent) under controlled conditions. Halogens add quantitatively
to the double bonds in the unsaturated fatty acids, principally oleic and linoleic
acids in the case of corn oil. Unreacted halogens are determined by titrating with
thiosulfate. The iodine number is defined as the grams of halogen, expressed as
iodine, reacting with 100 g of oil.

The concentration of standard Wijs solution, c(ICl) = 0.1 mol/L:

Hence its initial amount of substance n1(ICl) = c(ICl)*V(ICl) = 0.1*0.007 = 0.0007 mol

The amount of substance of the consumed for titration thiosulfate is equal to the residual amount of ICl:

n2(ICl) = c(Na2S2O3)*V(Na2S2O3) = 0.00124 mol

This amount is larger then the initial, so there is some mistake in the initial data. Please check it.

The final iodine number can be calculated in such way:

1) Obtain the mass of consumed iodine:
m1(I2) = n(I2)*M(I2) = 253.8*n(I2)

2) Obtain the mass of iodine necessary for 100 g of fat from proportion:
m(I2) = m1(I2)*100/0.11 = 909.1*m1(I2)